# Mean-field and fluctuations for hub dynamics in heterogeneous random networks

### v1

@@
Abstract. In a class of heterogeneous random networks, where each node dynamics is a random dynamical system, interacting with neighbor nodes via a random coupling function, we characterize the hub behavior as the mean-field, subject to statistically controlled fluctuations. In particular, we prove that the fluctuations are small over exponentially long time scales and obtain Berry-Esseen estimates for the fluctuation statistics at any fixed time. Our results provide a mathematical explanation for several numerical observations, including the scaling relation between system size and frequency of large fluctuations, as well as system size induced desynchronization. To our best knowledge, these are the first characterizations of mean-field fluctuations on networks with a degree distribution that follows a power-law, a common feature for many realistic systems.
@@ 1 Introduction
Network systems are fruitful models for various naturally occurring and man-made systems ranging from neuroscience [BGP+09] and physics [MS09] via electrochemistry [NOEE+22], to social sciences [SPMS17] to mention a few applications. In the case of homogeneous networks, where symmetries facilitate the analysis, many ergodic and statistical properties of the network system are known, in the context of coupled map lattices [CF05], all-to-all coupled systems [Fer19, BLS23], and in the all-to-all thermodynamic limit [ST21, ST22]. Much less is known about heterogeneous networks, another class of realistic models which generally lack symmetry and feature massively connected nodes, referred to as hubs, coexisting with poorly connected nodes [JSS13].
@@
In this paper, we develop new probability theory to meet these challenges. By exploiting the typicality of random trajectories, we overcome some major technical challenges arising in the gap between topological dynamics and ergodic theory. Furthermore, by concentration inequalities applied to the Chung-Lu random power-law network, we establish the locally star-like property of hubs. As a result, we advance the state of art of probability theory for network dynamics and complex systems by rigorously characterizing hub dynamics in random power-law networks, in terms of the mean-field subject to Gaussian-like fluctuations.
@@ 1.1 Network random dynamical system
$x_{i}^{t+1}=f_{\omega_{i}^{t}}(x_{i}^{t})+\frac{\alpha}{\Delta_{0}}\sum_{j=1}^{N}A_{ij}h_{\omega_{i}^{t}}(x_{i}^{t},x_{j}^{t})\mod 1,~{}~{}~{}~{}i=1,\cdots,N.$ (1)
@@
In this example the node dynamics admit $\mathrm{Leb}_{\mathbb{T}}$ as the unique stationary measure. The stationary measure can be absolutely continuous or singular with respect to $\mathrm{Leb}_{\mathbb{T}}$, including measures supported on a Cantor set; Theorem 3.3 will cover a more general case.
@@ 1.2 Dynamics on the star network
$z^{t+1}=$ $f_{\omega_{1}^{t}}(z^{t})+\frac{\alpha}{L}\sum_{j=2}^{N}h_{\omega_{1}^{t}}(z^{t},x_{j}^{t})\mod 1,$ (4)
$x_{j}^{t+1}=$ $f_{\omega_{j}^{t}}(x_{j}^{t})+\frac{\alpha}{L}h_{\omega_{j}^{t}}(x_{j}^{t},z^{t})\mod 1,~{}~{}~{}~{}j=2,\cdots,N.$
@@ 1.2.1 Mean-field dimensional reduction
We simulate the star network by probing three coupling strengths $\alpha=0.05$, $0.8$, $0.9$. For each $\alpha$, we initialize the node states $(z^{0},x_{2}^{0},\cdots,x_{N}^{0})\in\mathbb{T}^{N}$ at random with uniform distribution in $[0,1)$. Then, we iterate Eq. (4), discard the first $5000$ iterates, and collect the next $1000$ iterates. In Figure 2, we plot the hub return map $z^{t}$ versus $z^{t+1}$ in red. As the coupling strength $\alpha$ varies, the hub behavior differs from the isolated node dynamics. In Figure 1 on the left panel at $\alpha=0.05$, the hub remains contractive, on the central panel at $\alpha=0.8$, the hub dynamics appear to have an expanding region, and lastly, on the right panel at $\alpha=0.9$, the hub appears to hover around a fixed point near $0.2$. This shows the variety of hub behaviors emergent from the interactions, according to different coupling strengths.
@@
$\frac{1}{L}\sum_{j=2}^{N}h_{\omega_{1}^{t}}(z^{t},x_{j}^{t})$ $\approx\int_{\mathbb{T}}h_{\omega_{1}^{t}}(z^{t},x)\mathrm{d}x$
$=\sin 2\pi z^{t}-\frac{\omega_{1}^{t}}{3.6}$
@@
$z^{t+1}=f_{\alpha,\omega_{1}^{t}}(z^{t})+\xi^{t},$
@@
$f_{\alpha,\omega}(z):=f_{\omega}(z)+\alpha\int_{\mathbb{T}}h(\omega,z,x)\mathrm{d}x\mod 1,$ (5)
@@
$\xi^{t}(\omega,x):=\frac{\alpha}{L}\sum_{j=2}^{N}\sin 2\pi x_{j}^{t}.$
@@ 1.2.2 Frequency of large fluctuations
Figure 3: Frequency of large fluctuations decreases exponentially in system size. The red diamonds mark the frequency $\rho_{\varepsilon}^{T}$ up to time $T=2\times 10^{5}$ of large mean-field fluctuations, i.e., $|\xi^{t}|>\varepsilon$ with threshold $\varepsilon=0.025$. The horizontal axis for system size $L$ is in linear scale, whereas the vertical axis for frequency $\rho_{\varepsilon}^{T}$ is in logarithmic scale. The green line provides a tight linear fit, indicating exponential decrease of $\rho_{\varepsilon}^{T}$ in $L$.
@@ 1.2.3 Gaussian behavior of the fluctuations
The fluctuation $\xi^{t}$ can be interpreted as an ensemble average of the low degree node states through the observable $x\mapsto\alpha\sin 2\pi x$. The low degree nodes are almost isolated, up to hub influence of order $O(L^{-1})$, and hence almost independent from each other. Hence we expect the fluctuations $\xi^{t}$ to follow a Central Limit behavior. To illustrate this, we fix star size $L=10^{4}$ and coupling strength $\alpha=0.9$, and take $10^{4}$ trials of network initial conditions randomly independently and uniformly in $\mathbb{T}^{L+1}$. For each trial $n$ we simulate the star network dynamics up to time $T=1000$ and calculate the terminal fluctuation $\xi^{T}_{n}$. We plot the histogram of the data $\{\xi^{T}_{n}\}_{n=1}^{10^{4}}$ in Figure 4. Superimposed in green is the probability density function of the normal distribution $\mathcal{N}(0,\frac{\alpha^{2}}{2L})$ with zero mean and variance $\frac{\alpha^{2}}{2L}$. The close fit indicates that $\mathcal{N}(0,\frac{\alpha^{2}}{2L})$ indeed captures the fluctuation statistics at time $T$; see Theorem A (iii).
@@ 1.2.4 Statement of main result on the star network
Theorem A below is a particular version of the more general Theorem 2.3 and underpins the observations made for the star network dynamics (4): the mean-field dimensional reduction for hub behavior in Section 1.2.1 is addressed in item (i), scaling relations for the large fluctuation frequency in Section 1.2.2 are addressed in item (ii), and finally, the Gaussian nature of fluctuations in Section 1.2.3 are addressed in item (iii).
In the statement below, we use $\mathrm{Prob}$ to denote the product probability measure on $\Omega\times\mathbb{T}^{N}$ of the Bernoulli measure for $\Omega$ times the volume for $\mathbb{T}^{N}$.
Theorem A: hub dynamics in star network random dynamical system. Consider dynamics (4) on a star network with $L\gg 1$ and initial conditions following the uniform distribution on $\mathbb{T}^{N}$. Then, at coupling strength $\alpha\ll L^{1/2}$, the hub evolution admits mean-field dimensional reduction defined in (5), namely:
@@ 1. $\mathrm{(i)}$
    $\mathrm{Prob}\left\{(\omega,x):\liminf_{T\to+\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbbm{1}_{|\xi^{t}(\omega,x)|\leq\varepsilon}\geq 1-\exp(-L\varepsilon^{2}\alpha^{-2}/9)\right\}=1;$
@@ 2. $\mathrm{(ii)}$
    $\mathrm{Prob}\left\{(\omega,x):\max_{t\in I_{t_{0}}^{T}}|\xi^{t}(\omega,x)|\leq 3L^{-\kappa}\alpha\right\}\geq 1-\exp(-L^{1-2\kappa}),~{}~{}~{}~{}\kappa\in(0,1/2);$
@@ 3. $\mathrm{(iii)}$
    $\mathrm{Prob}\left\{(\omega,x):\xi^{t}(\omega,x)\leq s\right\}\in\left[F_{L}\left(s-c_{1}\right)-c_{2},F_{L}\left(s+c_{1}\right)+c_{2}\right],~{}~{}~{}~{}s\in\mathbb{R},$
@@
In Items (i) and (ii), we have provided an upper bound for the fluctuation size and a lower bound for the asymptotic frequency; the constants $A,\gamma$ in Figure 3 are not sharp and generally may depend on the coupling function. Theorem A is obtained as a special case of more general results, see Theorem 2.3. We briefly discuss the proof strategy in this particular case, which will become a fundamental step in Theorem 2.16 for dimensional reduction on more complex networks.
@@
Item (iii) follows from Berry-Esseen estimates together with our shadowing technique in step 3. We consider the isolated node dynamics, observed through $x\mapsto\sin 2\pi x$, as iid random variables on $\Omega\times\mathbb{T}^{L+1}$. The fluctuation $\xi^{t}$ is thus the ensemble average, whose Gaussian nature conforms to the classic Berry-Esseen estimates. Our result follows by comparing orbit-wise the isolated dynamics to the low degree trajectory as in step 3.
@@ 1.3 Dynamics on power-law networks
where $n=10^{6}$ denotes the number of nodes, $\beta=3$ power-law exponent , $w=10$ average expected degree, and $m=10^{3}$ largest expected degree. On the empty graph consisting of nodes $1,\cdots,n$, we add links between nodes $i$ and $j$ as Bernoulli variables with success probability $p_{ij}=w_{i}w_{j}/\sum_{k=1}^{n}w_{k}$. The resulting graph $G_{0}$ is not connected, but has a giant component, which we take to be $G_{1}$. By concentration inequalities, we will see in Lemma 3.6 that the actual degree sequence $k_{i}$ is concentrated at the expected version $w_{i}$ above; moreover, we will prove that this random power-law graph is locally star-like in the sense that most neighbors of any hub are of low degree, see Theorem B. In our simulation, $G_{1}$ consists of $N=998168$ nodes, with maximum degree $\Delta_{0}=979$ and minimum degree 1, and is a connected power-law graph. To showcase its power-law degree distribution, we plot in Figure 5 left panel the degree $k$ against the frequency $P(k)$ of nodes of degree $k$ in log-log scale. For comparison, we show a power-law $k^{-3}$ in green.
@@
Figure 5 left panel shows that there are no natural scales to distinguish the hubs from low degree nodes, so we have to introduce them by hand. For $G_{1}$, we put hub scale $\Delta=900$, low degree scale $\delta=100$, and thus define the collection of $\Delta$-hubs to be
@@
In $G_{1}$, we find $\nu=0.941$; in other words, more than $94.1\%$ of neighbors of each of the seven hubs in $G_{1}$ are of low degree. As we will prove in Theorem 3.7, the star-like index $\nu$ of any large power-law network $G$ is close to 1, given the appropriate scales $\Delta,\delta$.

### v2

@@
Abstract. We study a class of heterogeneous random networks, where the network degree distribution follows a power-law, and each node dynamics is a random dynamical system, interacting with neighboring nodes via a random coupling function. We characterize the hub behavior by the mean-field, subject to statistically controlled fluctuations. In particular, we prove that the fluctuations are small over exponentially long time scales and obtain Berry-Esseen estimates for the fluctuation statistics at any fixed time. Our results provide an explanation for several numerical observations, namely, a scaling relation between system size and frequency of large fluctuations, the system size induced desynchronization, and the Gaussian behavior of the fluctuations.
Data availability: We do not analyse or generate any datasets, because our work proceeds within a theoretical and mathematical approach.
All authors have no conflicts of interest.
@@ 1 Introduction
Network systems are fruitful models for various naturally occurring and man-made systems ranging from neuroscience [BGP+09] and physics [MS09] via electrochemistry [NOEE+22], to social sciences [SPMS17] to mention a few applications. In the case of homogeneous networks, where symmetries facilitate the analysis, many ergodic and statistical properties of the network system are known, in the context of coupled map lattices [CF05], all-to-all coupled systems [Fer19, BLS23], and in the all-to-all thermodynamic limit [ST21, ST22, Gal22]. Much less is known about heterogeneous networks, another class of realistic models which generally lack symmetry and feature massively connected nodes, referred to as hubs, coexisting with poorly connected nodes [JSS13].
@@
In this paper, we meet these challenges by exploiting the typicality of random trajectories, and thereby overcoming some major technical challenges arising in the gap between topological dynamics and ergodic theory. We also apply concentration inequalities to the Chung-Lu random power-law network to establish certain graph theoretic properties. In particular, we advance the state-of-the-art of ergodic theory for network dynamics and complex systems by characterizing hub dynamics in random power-law networks, in terms of the mean-field subject to Gaussian-like fluctuations. We exhibit examples of uniformly contracting node dynamics in Introduction and Section 3 and comment on expanding and nonuniformly contracting cases in Remarks 2.5, 3.4.
@@ 1.1 Network random dynamical system
$x_{i}^{t+1}=f_{\boldsymbol{\omega}_{i}^{t}}(x_{i}^{t})+\frac{\alpha}{\Delta_{0}}\sum_{j=1}^{N}A_{ij}h_{\boldsymbol{\omega}_{i}^{t}}(x_{i}^{t},x_{j}^{t})\mod 1,~{}~{}~{}~{}i=1,\cdots,N.$ (1)
@@
Remark 1.1 (Notation of noise realization).
In the introduction, we use the simple font $\omega$ to index the family of circle maps $f_{\omega}$ and coupling functions $h_{\omega}$, and the boldface $\boldsymbol{\omega}=(\boldsymbol{\omega}_{i}^{t})$ to represent a vector of noise realizations with time $t=0,1,\cdots$ and node $i=1,\cdots,N$ coordinates.
@@
In this example the node dynamics admit $\mathrm{Leb}_{\mathbb{T}}$ as the unique stationary measure. For other choices of $f_{\omega}$, the stationary measure can be absolutely continuous or singular with respect to $\mathrm{Leb}_{\mathbb{T}}$, including measures supported on a Cantor set; Theorem 3.3 will cover a more general case.
@@
In the rest of the Introduction, we continue with the concrete examples of node dynamics $f_{\omega}$ and coupling function $h_{\omega}$ as in Example 1.2.
@@ 1.2 Dynamics on the star network
$z^{t+1}=$ $f_{\boldsymbol{\omega}_{1}^{t}}(z^{t})+\frac{\alpha}{L}\sum_{j=2}^{N}h_{\boldsymbol{\omega}_{1}^{t}}(z^{t},x_{j}^{t})\mod 1,$ (4)
$x_{j}^{t+1}=$ $f_{\boldsymbol{\omega}_{j}^{t}}(x_{j}^{t})+\frac{\alpha}{L}h_{\boldsymbol{\omega}_{j}^{t}}(x_{j}^{t},z^{t})\mod 1,~{}~{}~{}~{}j=2,\cdots,N.$
Here the noise $\boldsymbol{\omega}_{i}^{t}$ is assumed to be iid in time $t=0,1,\cdots$ and in node coordinates $i=1,\cdots,N$, assigning weight $1/4$ to each $\omega\in\{0,1,2,3\}$.
@@ 1.2.1 Mean-field dimensional reduction
We simulate the star network by probing three coupling strengths $\alpha=0.05$, $0.8$, $0.9$. For each $\alpha$, we initialize the node states $(z^{0},x_{2}^{0},\cdots,x_{N}^{0})\in\mathbb{T}^{N}$ at random with uniform distribution in $[0,1)$. Then, we iterate Eq. (4), discard the first $5000$ iterates, and collect the next $1000$ iterates. In Figure 2, we plot the hub return map $z^{t}$ versus $z^{t+1}$ in red. As the coupling strength $\alpha$ varies, the hub behavior differs from the isolated node dynamics. In Figure 1 on the left panel at $\alpha=0.05$, the hub remains contractive, on the central panel at $\alpha=0.8$, the hub dynamics appear to have an expanding region, and lastly, on the right panel at $\alpha=0.9$, the hub appears to hover around a fixed point near $0.2$. This shows the variety of hub behaviors emergent from the interactions, according to different coupling strengths. Due to our choice of $h_{\omega}$ in Eq. (3), the randomness in $f_{\alpha,\omega}$ collapses at $\alpha=0.9$, resulting in the single graph of $f_{0.9}=f_{0.9,\omega}$, $\omega=0,1,2,3$ as shown in the right panel of Figure 2.
@@
$\frac{1}{L}\sum_{j=2}^{N}h_{\boldsymbol{\omega}_{1}^{t}}(z^{t},x_{j}^{t})$ $\approx\int_{\mathbb{T}}h_{\boldsymbol{\omega}_{1}^{t}}(z^{t},x)\mathrm{d}x$
$=\sin 2\pi z^{t}-\frac{\boldsymbol{\omega}_{1}^{t}}{3.6}$
@@
$z^{t+1}=f_{\alpha,\boldsymbol{\omega}_{1}^{t}}(z^{t})+\xi^{t}(\boldsymbol{\omega},x),$
@@
$f_{\alpha,\omega}(z):=f_{\omega}(z)+\alpha\int_{\mathbb{T}}h_{\omega}(z,x)\mathrm{d}x\mod 1,$ (5)
@@
$\xi^{t}(\boldsymbol{\omega},x):=\frac{\alpha}{L}\sum_{j=2}^{N}\sin 2\pi x_{j}^{t}.$
@@ 1.2.2 Frequency of large fluctuations
Figure 3: Frequency of large fluctuations decreases exponentially in system size. The red diamonds mark the frequency $\rho_{\varepsilon}^{T}$ up to time $T=2\times 10^{5}$ of large mean-field fluctuations, i.e., $|\xi^{t}|>\varepsilon$ with threshold $\varepsilon=0.025$. The horizontal axis for system size $L$ is in linear scale, whereas the vertical axis for frequency $\rho_{\varepsilon}^{T}$ is in logarithmic scale. The green line provides a tight linear fit, indicating an exponential decrease of $\rho_{\varepsilon}^{T}$ in $L$.
@@ 1.2.3 Gaussian behavior of the fluctuations
The fluctuation $\xi^{t}$ can be interpreted as an ensemble average of the low degree node states through the observable $x\mapsto\alpha\sin 2\pi x$. The low degree nodes are almost isolated, up to hub influence of order $O(L^{-1})$, and hence almost independent from each other. Hence, we expect the fluctuations $\xi^{t}$ to follow a Central Limit behavior. To illustrate this, we fix star size $L=10^{4}$ and coupling strength $\alpha=0.9$, and take $10^{4}$ trials of network initial conditions randomly independently and uniformly in $\mathbb{T}^{L+1}$. For each trial $n$ we simulate the star network dynamics up to time $T=1000$ and calculate the terminal fluctuation $\xi^{T}_{n}$. We plot the histogram of the data $\{\xi^{T}_{n}\}_{n=1}^{10^{4}}$ in Figure 4. Superimposed in green is the probability density function of the normal distribution $\mathcal{N}(0,\frac{\alpha^{2}}{2L})$ with zero mean and variance $\frac{\alpha^{2}}{2L}$. The close fit indicates that $\mathcal{N}(0,\frac{\alpha^{2}}{2L})$ indeed captures the fluctuation statistics at time $T$; see Theorem A (iii).
@@ 1.2.4 Statement of main result on the star network
Theorem A below underpins the observations made for the star network dynamics (4): the mean-field dimensional reduction for hub behavior in Section 1.2.1 is addressed in item (i), scaling relations for the large fluctuation frequency in Section 1.2.2 are addressed in item (ii), and finally, the Gaussian nature of fluctuations in Section 1.2.3 are addressed in item (iii).
In the statement below, we use $\mathrm{Prob}$ to denote the product probability measure on $\Omega\times\mathbb{T}^{N}$ of the Bernoulli measure for $\boldsymbol{\omega}\in\Omega=\{0,1,2,3\}^{\mathbb{N}\times N}$ times the volume for $x\in\mathbb{T}^{N}$.
Theorem A: hub dynamics in star network random dynamical system. Consider the dynamics (4) on a star network with $f_{\omega},h_{\omega}$ as in Example 1.2, and initial conditions following the uniform distribution on $\mathbb{T}^{N}$, $N=L+1\gg 1$. Then, at coupling strength $\alpha\ll L^{1/2}$, the hub evolution admits mean-field dimensional reduction defined in (5), namely:
@@ 1. $\mathrm{(i)}$
    $\mathrm{Prob}\left\{(\boldsymbol{\omega},x):\liminf_{T\to+\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbbm{1}_{|\xi^{t}(\boldsymbol{\omega},x)|\leq\varepsilon}\geq 1-\exp(-L\varepsilon^{2}\alpha^{-2}/9)\right\}=1;$
@@ 2. $\mathrm{(ii)}$
    $\mathrm{Prob}\left\{(\boldsymbol{\omega},x):\max_{t\in I_{t_{0}}^{T}}|\xi^{t}(\boldsymbol{\omega},x)|\leq 3L^{-\kappa}\alpha\right\}\geq 1-\exp(-L^{1-2\kappa}),~{}~{}~{}~{}\kappa\in(0,1/2);$
@@ 3. $\mathrm{(iii)}$
    $\mathrm{Prob}\left\{(\boldsymbol{\omega},x):\xi^{t}(\boldsymbol{\omega},x)\leq s\right\}\in\left[F_{L}\left(s-c_{1}\right)-c_{2},F_{L}\left(s+c_{1}\right)+c_{2}\right],~{}~{}~{}~{}s\in\mathbb{R},$
@@
Remark 1.3.
To state the relations among $L,\alpha,\varepsilon$ more precisely in Theorem A, we mean that there are constants $C_{1},C_{2},C_{3}>0$ such that if $L\geq C_{1}$, $L^{1/2}/\alpha\geq C_{2}$ and $\varepsilon L^{1/2}/\alpha\geq C_{3}$, then Items (i), (ii) and (iii) hold.
In Item (i), $\mathbbm{1}_{|\xi^{t}(\boldsymbol{\omega},x)|\leq\varepsilon}$ indicates whether or not the fluctuaion $\xi^{t}(\boldsymbol{\omega},x)$ at time $t$ is small; the time average computes the relative frequency of small fluctuations in the time window $t=0,\cdots,T-1$; by passing to the limit inferior we obtain the asymptotic frequency of small fluctuations starting from initial data $(\boldsymbol{\omega},x)$; finally, Item (i) says that with full probability, the asymptotic frequency of small fluctuations is close to one.
In Items (i) and (ii), we have provided an upper bound for the fluctuation size and a lower bound for the asymptotic frequency; the constants $A,\gamma$ in Figure 3 are not sharp and generally may depend on the coupling function. Theorem A is a consequence of the more general Theorem 2.4 and the uniform typicality of the random orbits of the contractions (2), see Theorem 3.3. The derivation of the explicit constants is discussed in Appendix A.1. We briefly discuss the proof strategy in this particular case, which will become a fundamental step in Theorem 2.18 for dimensional reduction on more complex networks.
@@
Item (iii) follows from Berry-Esseen estimates together with our shadowing technique in step 3. We consider the isolated node dynamics, observed through $x\mapsto\sin 2\pi x$, as iid random variables on $\Omega\times\mathbb{T}^{N}$. The fluctuation $\xi^{t}$ is thus the ensemble average, whose Gaussian nature conforms to the classic Berry-Esseen estimates. Our result follows by comparing orbit-wise the isolated dynamics to the low degree trajectory as in step 3.
@@ 1.3 Dynamics on power-law networks
where $n=10^{6}$ denotes the number of nodes, $\beta=3$ power-law exponent , $w=10$ average expected degree, and $m=10^{3}$ largest expected degree. On the empty graph consisting of nodes $1,\cdots,n$, we add links between nodes $i$ and $j$ as independent Bernoulli variables with success probability $p_{ij}=w_{i}w_{j}/\sum_{k=1}^{n}w_{k}$. The resulting graph $G_{0}$ is not connected, but has a giant component, which we take to be $G_{1}$. By concentration inequalities, we will see in Lemma 3.6 that the actual degree sequence $k_{i}$ is concentrated at the expected version $w_{i}$ above; moreover, we will prove that this random power-law graph is locally star-like in the sense that most neighbors of any hub are of low degree, see Theorem B. In our simulation, $G_{1}$ consists of $N=998168$ nodes, with maximum degree $\Delta_{0}=979$ and minimum degree 1, and is a connected power-law graph. To showcase its power-law degree distribution, we plot in Figure 5 left panel the degree $k$ against the frequency $P(k)$ of nodes of degree $k$ in log-log scale. For comparison, we show a power-law $k^{-3}$ in green.
@@
Figure 5 left panel shows no gap in the degree distribution; in particular, there are no natural scales to distinguish the hubs from low degree nodes, so we have to introduce them by hand. For $G_{1}$, we put hub scale $\Delta=900$, low degree scale $\delta=100$, and thus define the collection of $\Delta$-hubs to be
@@
In $G_{1}$, we find $\nu=0.941$; in other words, more than $94.1\%$ of neighbors of each of the seven hubs in $G_{1}$ are of low degree. As we will prove in Theorem 3.7, the star-like index $\nu$ of a large power-law network $G$ with exponent $\beta>2$ is close to 1, given the appropriate scales $\Delta,\delta$.